The Linux Cyrillic HOWTO

Alexander L. Belikoff v2.1, 17 September 1995

This document describes how to set up your Linux box to typeset, view and print the documents in the Russian language.

Contents

1 Introduction 2

2 Acknowledgments and copyrights 2

3 Further plans 3

4 Setting up the environment 3 4.1 Console . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 3 4.2 The X Window System . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4

4.2.1 The X fonts. . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 4 4.2.2 The input translation . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 5

5 Cyrillic support in LaTeX 10 5.1 Using the cmcyralt package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 11 5.2 Using the CyrTUG package . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12

6 Miscellaneous utilities setup 12 6.1 bash . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 6.2 csh/tcsh . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 12 6.3 emacs . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 6.4 ispell . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 6.5 less . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 6.6 Netscape . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 13 6.7 rlogin . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

7 Printing 14 8 Useful Tools 14

8.1 User’s tools . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

8.2 Programmer’s tools . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . . 14

9 Summary of the various useful resources 15

1 Introduction

[ 27 June 1996 As far as I am concerned, this HOWTO is not being maintained by the author any more. If you are interested in maintaining the Cyrillic HOWTO, or know how to contact the author, please get in touch with me at gregh@sunsite.unc.edu ]

This document covers the things you need to successfully typeset, view, and print documents in Russian.

The description is targeted primarily to the end users. While writing it, I tried to keep the things as simple as possible. On my opinion, it is unfair to overburden the user with the X Window System or TeX implementation details, because she merely wants to type or view the nice texts in Russian. However, I don’t describe what the X Window System is and how to typeset the documents with TeX and LaTeX. Consequently, having a guru at hand for the case of potential problems would be an advantage.

There is some conflict between MS-DOG and Un*x Cyrillic codesets. In MS-DOG, the most popular Cyrillic codeset is Alt (so-called alternative codeset). In Un*x, however, the traditional codeset with Russian characters is KOI-8. It is specified in the RFC 1489 (”Registration of a Cyrillic Character Set”).

The difference in two codesets is usage is pretty minimal (except the TeX packages). Therefore, we will describe only KOI-8 codeset setup.

I appreciate any comments corrections and suggestions concerning the document.

2 Acknowledgments and copyrights

Many people helped me (and not only me) with valuable information and suggestions. Even more people contributed software to the public community. I am sorry if I have forgotten to mention somebody.

So, here they go: Bas V. de Bakker, David Daves, Serge Vakulenko, Sergei O. Naoumov, Winfried Truemper.

This document is Copyright (C) 1995 by Alexander L. Belikoff. It may be used and distributed under the usual Linux HOWTO terms described below.

The following is a Linux HOWTO copyright notice:

Unless otherwise stated, Linux HOWTO documents are copyrighted by their respective authors. Linux HOWTO documents may be reproduced and distributed in whole or in part, in any medium physical or electronic, as long as this copyright notice is retained on all copies. Commercial redistribution is allowed and encouraged; however, the author would like to be notified of any such distributions.

All translations, derivative works, or aggregate works incorporating any Linux HOWTO documents must be covered under this copyright notice. That is, you may not produce a derivative work from a HOWTO and impose additional restrictions on its distribution. Exceptions to these rules may be granted under certain conditions; please contact the Linux HOWTO coordinator at the address given below.

In short, we wish to promote dissemination of this information through as many channels as possible. However, we do wish to retain copyright on the HOWTO documents, and would like to be notified of any plans to redistribute the HOWTOs.

If you have questions, please contact Greg Hankins, the Linux HOWTO coordinator, at gregh@sunsite.unc.edu. You may finger this address for phone number and additional contact information.

3 Further plans

The next versions of this document will be accessible on sunsite.unc.edu and tsx-11.mit.edu in HOWTO directory of the Linux Document Project.

I am looking forward to include more information on TeX and LaTeX as well as much more elaborated information on printing.

4 Setting up the environment

4.1 Console

All stuff needed for russification of the Linux console is contained in the kbd package. The package is accessible at sunsite.unc.edu or tsx-11.mit.edu. Usually, that package is already installed (it is a standard part of at least Slackware distribution).

To setup the Cyrillic stuff, one should do three things:

1. Set the appropriate screen font. This is performed by the setfont program. The fonts files are placed

in /usr/lib/kbd/consolefonts.

NOTE: Never run the setfont program under X or it will hang your system. This is because it works with low-level video card calls which X doesn’t like.

2. If you use the font in Alt coding (as I do) then you have to set up the screen mapping program to perform automatic conversion from Alt to KOI-8. For that purpose use the mapscrn program and the

/usr/lib/kbd/consoletrans/koi2alt file.

3. Load the appropriate keyboard layout with the loadkeys program.

4. Output an ESC (K escape sequence on the screen (ESC stands for the Escape character with code 033). Only God knows the purpose of that combination. I stole it from the Danish-HOWTO

     <file://sunsite.unc.edu/pub/Linux/docs/HOWTO/Danish-HOWTO> (thanks, Thomas Petersen) and

it works for me!

The following is an example of a script which sets up the Cyrillic mode for console:

#!/bin/bash
#
# load cyrillic defs for console
#
# *** NEVER TRY IT UNDER X!!! ***
loadkeys /usr/lib/kbd/keytables/ru.map
setfont /usr/lib/kbd/consolefonts/Cyr_a8x16
mapscrn /usr/lib/kbd/consoletrans/koi2alt
echo -ne 'ESC(K'        # use the REAL ESCAPE character here !
echo "Use the right Ctrl key to switch the mode..."

4.2 The X Window System

Like the console mode, the X environment also requires some setup. This involves setting up the input mode and the X fonts. Both are being discussed below.

4.2.1 The X fonts.

First of all, you have to obtain the fonts collection having the Cyrillic glyphs at the appropriate places.

There is a number of such fonts on the net. The author’s favorite one is the collection VakuFonts created by Serge Vakulenko (vak@cronyx.ru). It can be found in the collection of cyrillic stuff for the X Window System

<ftp://ftp.funet.fi/pub/culture/russian/comp/xwin> where you can find many useful packages for X.

Note: Apparently, that package is included to the XFree86 version 3.1.2 as well as to the most recent public patch for the X11 Release 6. Unfortunately, the author hasn’t had a chance to check it yet.

Usually the X fonts are distributed in the BDF format which is actually the textual font description. You should compile the fonts to the PCF format using the bdftopcf command:

     bdftopcf -o name.pcf name.bdf

It is also possible to compress the compiled font using the compress program (I am not sure about the gzip support).

Now you should do three things to set everything up:

1. Put the compiled (and possibly compressed) fonts to the specified directory.

2. Recreate the list of fonts for the directory. Simply cd to it and run:

          mkfontdir .

You should run it once. This will upgrade the the fonts catalog file fonts.dir.

3. If the fonts package provides the file of fonts’ aliases (usually fonts.alias) then append it’s contents to the fonts’ aliases file in the directory containing the fonts.

4. If that directory is not already known to the X server, then you should make it known. To achieve that, add the following commands to the xinitrc file (either local or global one):

          xset +fp directory with fonts xset fp rehash

After you have made the settings above, you can check the availability of the new fonts by running the following command:

     xfd -fn fontname

This should show the table of characters of the specified font.

4.2.2 The input translation

The switching between the different input translations is set up by the xmodmap program. This program allows customization of codes emitted by various characters and their combinations. It sets the things up based on the file containing the translation table, usually ~/.Xmodmap.

The following is a simplified description of input customization. If you want to do more sophisticated tricks, refer to the xmodmap(1) or, even better, wait for the X11 Release 7 which will address the current input problems.

In our case, the translation table should define two things:

• the character codes emitted by the alphanumeric characters, and

• the mode switching rules

The table of characters This is basically a sequence of directives which assign the certain keysyms to a specified keycodes. The general syntax is the following:

keycode code = sym1 sym2 sym3 sym4

where code is the numerical code of the given key on the keyboard (refer to the standard table for your system. In my case it is stored in the file /usr/lib/X11/etc/xmodmap.std). The syms define the keysyms emitted by that key in different conditions. Sym1 is the keysym emitted by the key in a regular state, sym2 corresponds the key in shifted state (usually when Shift is held down). Sym3 and sym4 define the keysyms emitted when the Mode switch is active for the normal and shifted states respectively (group 2 , according the X Protocol Specification). In our case, the active Mode switch corresponds to the Cyrillic input mode.

These should be either hexadecimal codes or the symbolic constants from /usr/include/X11/keysymdef.h (without leading ”XK ”).

Thus, if we wanted the key corresponding to the Latin ’a’ generate the Russian ’a’ in the alternative mode, we would write the following:

keycode 38  =  a  A  0xC1 0xE1

The reader might be curious why I haven’t used the Cyrillic a and Cyrillic A constants respectively. The answer is that it didn’t work for me. I am not very familiar with the guts of the X Window System specification, but I have the following explanation. The symbolic constants above have the values 0x6C1 and 0x6E1 respectively. This means that in really multi-lingual environment they could be successfully used without overlapping with any other character set. However the KOI-8 standard is not well suited for such environment. Thus, since we want to retain compatible with the past, we will violate the rules of multi-lingual support in the X Window System.

The following is a table for the most popular russian JCUKEN keyboard layout (these tables are derived from the ones in the VakuFonts package):

keysym  4             = 4               dollar          4       quotedbl
keysym  5             = 5               percent         5       colon
keysym  6             = 6               asciicircum     6       comma
keysym  7             = 7               ampersand       7       period
keysym  q             = q               Q               0xCA    0xEA
keysym  w             = w               W               0xC3    0xE3
keysym  e             = e               E               0xD5    0xF5
keysym  r             = r               R               0xCB    0xEB
keysym  t             = t               T               0xC5    0xE5
keysym  y             = y               Y               0xCE    0xEE
keysym  u             = u               U               0xC7    0xE7
keysym  i             = i               I               0xDB    0xFB
keysym  o             = o               O               0xDD    0xFD
keysym  p             = p               P               0xDA    0xFA
keysym  bracketleft   = bracketleft     braceleft       0xC8    0xE8
keysym  bracketright  = bracketright    braceright      0xDF    0xFF
keysym  a             = a               A               0xC6    0xE6
keysym  s             = s               S               0xD9    0xF9
keysym  d             = d               D               0xD7    0xF7
keysym  f             = f               F               0xC1    0xE1
keysym  g             = g               G               0xD0    0xF0
keysym  h             = h               H               0xD2    0xF2
keysym  j             = j               J               0xCF    0xEF
keysym  k             = k               K               0xCC    0xEC

keysym  l             = l               L               0xC4    0xE4
keysym  semicolon     = semicolon       colon           0xD6    0xF6
keysym  apostrophe    = apostrophe      quotedbl        0xDC    0xFC
keysym  grave         = grave           asciitilde      0xA3    0xB3
keysym  z             = z               Z               0xD1    0xF1
keysym  x             = x               X               0xDE    0xFE
keysym  c             = c               C               0xD3    0xF3
keysym  v             = v               V               0xCD    0xED
keysym  b             = b               B               0xC9    0xE9
keysym  n             = n               N               0xD4    0xF4
keysym  m             = m               M               0xD8    0xF8
keysym  comma         = comma           less            0xC2    0xE2
keysym  period        = period          greater         0xC0    0xE0

Also, for those using the russian YAWERTY layout, I’ve included the following table:

keysym  q             = q               Q               0xD1    0xF1
keysym  w             = w               W               0xD7    0xF7
keysym  e             = e               E               0xC5    0xE5
keysym  r             = r               R               0xD2    0xF2
keysym  t             = t               T               0xD4    0xF4
keysym  y             = y               Y               0xD9    0xF9
keysym  u             = u               U               0xD5    0xF5
keysym  i             = i               I               0xC9    0xE9
keysym  o             = o               O               0xCF    0xEF
keysym  p             = p               P               0xD0    0xF0
keysym  bracketleft   = bracketleft     braceleft       0xDB    0xFB
keysym  bracketright  = bracketright    braceright      0xDD    0xFD
keysym  a             = a               A               0xC1    0xE1
keysym  s             = s               S               0xD3    0xF3
keysym  d             = d               D               0xC4    0xE4
keysym  f             = f               F               0xC6    0xE6
keysym  g             = g               G               0xC7    0xE7
keysym  h             = h               H               0xC8    0xE8
keysym  j             = j               J               0xCA    0xEA
keysym  k             = k               K               0xCB    0xEB
keysym  l             = l               L               0xCC    0xEC
keysym  z             = z               Z               0xDA    0xFA
keysym  x             = x               X               0xD8    0xF8
keysym  c             = c               C               0xC3    0xE3
keysym  v             = v               V               0xD6    0xF6
keysym  b             = b               B               0xC2    0xE2
keysym  n             = n               N               0xCE    0xEE
keysym  m             = m               M               0xCD    0xED

keysym  backslash     = backslash       bar             0xDC    0xFC
keysym  grave         = grave           asciitilde      0xC0    0xE0
keysym  equal         = equal           plus            0xDE    0xFE
keysym  3             = 3               numbersign      3       0xDF
keysym  4             = 4               dollar          4       0xFF

The mode switching rules This is basically the trickiest part of the X Cyrillic setup. You should define the conditions in which the current mode is switched between the regular and the Cyrillic one.

There are two ways to achieve that in Linux. One is XFree86-specific, while the other is more general (well, not too much, as I’ll show below).

The XFree86-specific way is the following. There are two virtual actions which can be assigned to the keys in the XF86Config file: ModeShift which changes to the mode alternative to the regular one without locking, and ModeLock which does the same but with locking. In the first case the keys will emit the alternative keysyms only when the key generating the ModeShift is held down, whereas in the latter case the user needs to press the key generating the ModeLock keysym only once and the keyboard will be generating the alternative keysyms until that key is pressed for a second time. You should assign the ModeShift and ModeLock keysyms to the keys you want to work the mode switches.

Thus, if one wants to assign the ModeShift action to the right Alt key, she should place the following directive in her XF86Config:

RightAlt        ModeShift

Similarly, if the action required was ModeLock, the directive would be:

RightAlt        ModeLock

See the XF86Config(4/5) for more details.

The other way is, again, to use the xmodmap utility. This is much more tricky. Basically what you should do is:

• Assign the Mode switch keysym to some key, and

• Add Mode switch to some spare modifier map

Now the key to which the ModeShift is assigned will act as a mode switch. This means that while it is held down, the keyboard is in alternative mode.

Moreover, if you add a lockable key to that modifier’s map, this key will lock the alternative mode.

Note: There are some problems however. Serge Vakulenko (vak@cronyx.com) pointed out that the different X Server implementations may have different rules of assignments the mode switches (like, for example, some servers restrict the set of the keys which may work in toggle mode to, say, CapsLock, NumLock, and ScrollLock). Hopefully, this is a subject to change in the next release of the X Window System. For more details, see the X Protocol specification.

Let’s see an example. Suppose, one wants to use the right Alt as a mode switch and the ScrollLock as as a mode lock. First of all, one should check the default modifiers’ map. This is accomplished by running the xmodmap without arguments:

     $ xmodmap
     xmodmap:  up to 2 keys per modifier, (keycodes in parentheses):
     shift       Shift_L (0x32),  Shift_R (0x3e)
     lock        Caps_Lock (0x42)
     control     Control_L (0x25)
     mod1        Alt_L (0x40),  Alt_R (0x71)
     mod2        Num_Lock (0x4d)
     mod3
     mod4
     mod5

According to the above, the plan of attack is the following:

1. remove the Alt R key from the mod1 map

2. assign the Mode switch keysym to the Alt R key

3. assign the Scroll Lock keysym to the keycode 78 (the code of the actual ScrollLock)

4. add the Mode switch to the spare (mod3) map, and

5. add the Scroll Lock keysym to the mod3 map

Thus, here is the solution:

remove mod1 = Alt_R
keysym Alt_R = Mode_switch
keycode 78 = Scroll_Lock
add mod3 = Mode_switch
add mod3 = Scroll_Lock

If you use the latter solution, you may combine both the table and the mode directives in your ~/.Xmodmap file.

Such files are generally supplied with the various X Cyrillic stuff packages. The good example is the tables in the perfect package by Serge Vakulenko described above.

Once you have such file containing the table, you should run the command:

     xmodmap filename

every time you start X. Modify your .xinitrc file to perform it. NOTE: your .xinitrc can already contain the code to run the xmodmap over your local table if the one exists.

The table distributed with the Serge’s Vakulenko package didn’t work for the author. The following patch fixed the problem:

diff -u --new-file jcuken.xmm jcuken.xmm.mod
--- jcuken.xmm  Mon May 20 09:11:36 1991
+++ jcuken.xmm.mod      Sun Aug 13 15:44:06 1995
@@ -2,6 +2,8 @@
 ! Cyrillic keyboard mapping table.
 ! Produced by Serge Vakulenko, <vak@kiae.su>, Moscow.
 !
+! Modified by Alexander L. Belikoff (abel@wisdom.weizmann.ac.il), 1995
+!
 ! Russian JCUKENG keyboard layout implemented.
 ! Cyrillic characters are entered in koi8 encoding.
 !
@@ -10,7 +12,9 @@
 ! Use CapsLock as rus/lat switch key.
 remove lock = Caps_Lock
-add mod2 = Caps_Lock
+keysym Caps_Lock = Mode_switch
+add mod2 = Mode_switch
+add lock = Mode_switch
 !       Key             Base            Shift           Caps    ShiftCaps
 !------------------------------------------------------------------------

This allowed me to use the Caps Lock key to switch between normal and Cyrillic input modes.

5 Cyrillic support in LaTeX

In this section I’ll describe the procedure of making LaTeX typeset Russian. LaTeX is a macro package for TeX offering the user many useful styles, templates, and commands. If you are making your mind on what package to use, I personally suggest LaTeX. Moreover, there are two versions of LaTeX available - 2.09 is the old one, while 2e is a new pre-3.0 release. If you are using LaTeX 2.09, then switch quickly to the 2e. The latter retains compatibility with the old one, but has much more features. Hopefully, version 3 will be released soon. So far I describe the LaTeX 2e setup.

I have an experience with two packages. One is the cmcyralt package by Vadim V. Zhytnikov

(vvzhy@phy.ncu.edu.tw) and Alexander Harin (harin@lourie.und.ac.za), and the other is the LH

package by the CyrTUG group with styles and hyphenation for LaTeX2e by Sergei O. Naoumov (serge@astro.unc.edu). I’ll describe both.

Note: Both of these packages require the Cyrillic text to be typeset using the Alt codeset, not KOI-8 ! This is caused by historical reasons, since the creators of these packages used to work with EmTeX - the MS-DOG version of TeX (they didn’t know about Linux yet :-). Switching to the KOI-8 requires some effort and is being expected to be done soon. So far, use some utility to convert your russian text from KOI-8 to Alt. See section 8.1.

5.1 Using the cmcyralt package

The cmcyralt package can be found on any CTAN (Comprehensive TeX Archive Network) site like ftp.dante.de. You should get two collections: the fonts collection from fonts/cmcyralt and the styles

and hyphenation rules from macros/latex/contrib/others/cmcyralt.

Note: Make sure you have the Sauter package installed, since cmcyralt requires some fonts from it. You can get this package from CTAN site as well.

Now you should do the following:

1. Put the new fonts to the TeX fonts tree. On my system (Slackware 2.2) I created a cmcyralt directory in the /usr/lib/texmf/fonts/cm/. Create the src, tfm, and vf subdirectories in it. Put there .mf, .tfm, and vf files respectively.

2. Put the font driver files (*.fd) from the styles archive to the appropriate place (in my case it was

     /usr/lib/texmf/tex/latex/fd).

3. Put the style files (*.sty) to the appropriate LaTeX styles directory (in my case

     /usr/lib/texmf/tex/latex/sty).

Now the hyphenation setup. This requires to remake the LaTeX base file.

1. The file hyphen.cfg contains the directives for both English and Russian hyphenation. Extract the one for Russian and place it to the LaTeX hyphenation config file lthyphen.ltx. In my case, that file

     was in /usr/lib/texmf/tex/latex/latex-base.

2. Put the rhyphen.tex to the same directory. It is needed for making the new base file. Later, you can remove it.

3. Do ’make’ in that directory. Don’t for get to make a link from Makefile to Makefile.unx. During the make process check the output. There should be a message:

     Loading hyphenation patterns for Russian.

If everything goes OK, you will get the new latex.fmt in that directory. Put it to the appropriate place, where the previous one was (like /usr/lib/texmf/ini/). Don’t forget to save the previous one!.

This is it. The installation is complete. Try processing the examples found in the styles archive. If you are to create the PostScript files without any problems, then everything is OK. Now, to use Cyrillic in LaTeX, prepend your document with the following directive:

\usepackage{cmcyralt}

For more details, see the README file in the cmcyralt styles archive.

Note: if you do have problems with the examples, provided you have installed the things right, then probably your TeX system hasn’t been installed correctly. For example, during my first try, every attempt to create the .pk files for the russian fonts failed (MakeTeXPK stage). A substantial investigation discovered some implicit conflict between the localfont and ljfour METAFONT configurations. It used to work before, but kept crashing after the cmcyralt installation. Contact your local TeX guru - TeX is very (sometimes too much) complicated to reconfigure it without any prior knowledge.

5.2 Using the CyrTUG package

You  can  obtain  the  CyrTUG  package  from

the SunSite archive <ftp://sunsite.unc.edu/pub/academic/russian-studies/Software>. Get the files
CyrTUGfonts.tar.gz, CyrTUGmacro.tar.gz, and hyphen.tar.Z.

The process of installation doesn’t differ from the previous one.

6 Miscellaneous utilities setup

Generally, to set the certain utility up to handle the Cyrillic requires just to allow the 8 bit input. In some cases it is required to tell the application to show the extended ASCII characters in their ”native” form.

6.1 bash

Three variables should be set on order to make bash understand the 8-bit characters. The best place is ~/.inputrc file. The following should be set:

set meta-flag on
set convert-meta off
set output-meta on

6.2 csh/tcsh

The following should be set in .cshrc:

setenv LC_CTYPE iso_8859_5
stty pass8

If you don’t have the POSIX stty (impossible for Linux), then replace the last call to the following:

stty -istrip cs8

6.3 emacs

The minimal cyrillic support in emacs is done by adding the following calls to one’s .emacs (provided that the Cyrillic character set support is installed for console or X respectively):

(standard-display-european t)
(set-input-mode (car (current-input-mode))
   (nth 1 (current-input-mode))
   0)

This allows the user to view and input documents in Russian.

However, such mode is not of a big convenience because emacs doesn’t recognize the usual keyboard commands while set in Cyrillic input mode. There are a number of packages which use the different approach. They don’t rely on the input mode stuff established by the environment (either X or console. Instead, they allow the user to switch the input mode by the special emacs command and emacs itself is responsible for re-mapping the character set. The author took a chance to look at three of them. The russian.el package by Valery Alexeev (ava@math.jhu.edu) allows the user to switch between cyrillic and regular input mode and to translate the contents of a buffer from one Cyrillic coding standard to another (which is especially useful while reading the texts imported from MS-DOG). The rustable.el (sorry, I don’t know the author of it) adds the syntax rules of Cyrillic codeset to emacs (words’ bounds, case change rules etc.) These packages can be found at most Emacs-Lisp archives. Another one is the package remap which tries to make such support more generic. This package is written by Per Abrahamsen (abraham@iesd.auc.dk) and is accessible

at ftp.iesd.auc.dk.

As for the author’s opinion, I would suggest to start using the russian.el package because it is very easy to setup and use.

6.4 ispell

Check the sunsite.unc.edu:/pub/academic/russian-studies/Software for the russian dictionary cre-

ated by Neal Dalton (nrd@cray.com) for the ispell package.

6.5 less

So far, less doesn’t support the KOI-8 character set, but the following environment variable will do the job:

LESSCHARSET=latin1

6.6 Netscape

Set the following resource:

*documentFonts*registry:        koi8

6.7 rlogin

Use ’rlogin -8’

7 Printing

To print the text files containing the Russian characters using on PostScript printers, you need two things: the fonts and the 8bit-aware software able to print the texts using those fonts. The best package to print the text files in PostScript is a2ps by Evan Kirshenbaum (evan@csli) and Miguel Santana (miguel@imag.imag.fr). The last version is 8bit-aware. You can get it from imag.imag.fr:/archive/postscript/.

Check the sunsite.unc.edu:/pub/academic/russian-studies/Software/ in order to obtain the Cyril-

lic PostScript fonts. Also, there is a lot of fonts in the SimTel <ftp://oak.oakland.edu/pub/SimTel/> collection.

8 Useful Tools

8.1 User’s tools

There are number of programs able to convert from KOI-8 to Alt and back. You can even use the special mode for emacs (see section 6.3).

One  nice  standalone  package  is  translit.  It  is

available at the SunSite archive <ftp://sunsite.unc.edu/pub/academic/russian-studies/Software>.

This package is capable of converting between different formats including the KOI-8 and the Alt ones.

8.2 Programmer’s tools

So far, I explained the ways to make the programs accept and display the Cyrillic codeset. However the full localization of the system comprises much more. All discussed above is not enough. The system should be friendly for a user who doesn’t necessarily speak English. In my own opinion, it is not a big deal to become familiar with English at the level of the programs’ messages. However, it is not quite fair to require it.

Thus, the next level of localization requires the programs to be customizable to the requirements of different languages and data representation habits.

Before, that was done by developing some abstraction of the messages to output from the program’s code. Now, such mechanism is (more or less) standardized. And, of course, there are free implementations of it!

The good news is that GNU finally adopted the way of making the internationalized applications. Ulrich Drepper (drepper@ipd.info.uni-karlsruhe.de) developed a package gettext. This package is available at all GNU sites like prep.ai.mit.edu <ftp://prep.ai.mit.edu/pub/gnu/>. It allows you to develop programs in the way that you can easily make them support more languages. I don’t intend to describe the programming techniques, especially because the gettext package is delivered with excellent manual.

So, if you are developing programs which output messages (have you ever developed any program which didn’t?), then don’t be lazy to put a little (yes, really little) effort to make your program locale-aware.

Request for collaboration: If you want to learn the gettext package and to contribute to the GNU project simultaneously; or even if you just want to contribute, then you can do it! GNU goes international, so all the utilities are being made locale-aware. The problem is to translate the messages from English to Russian (and other languages if you’d like). Basically, what one has to do is to get the special .po file consisting of the English messages for a certain utility and to append each message with it’s equivalent in Russian. Ultimately, this will make the system speak Russian if the user wants it! For more details and further directions contact Ulrich Drepper (drepper@ipd.info.uni-karlsruhe.de <mailto:drepper@ipd.info.uni-karlsruhe.de>).

9 Summary of the various useful resources

The remap package for Emacs <ftp://ftp.iesd.auc.dk/>
Many fonts collections for X <ftp://ftp.funet.fi/pub/culture/russian/comp/xwin>
Information on Cyrillic Software <http://sunsite.oit.unc.edu/sergei/Software/Software.html>
Useful Cyrillic packages <ftp://sunsite.unc.edu/pub/academic/russian-studies/Software>
The kbd package for Linux <ftp://sunsite.unc.edu/pub/Linux/system/Keyboards/>
X fonts collections <ftp://ftp.switch.ch/mirror/linux/X11/fonts/>
rspell <ftp://sunsite.unc.edu/pub/academic/russian-studies/Software/rspell.tar.gz>